OpenStudy (anonymous):

Convert 15/50 to a percent. 0.3% 3.0% 30%

OpenStudy (anonymous):

This should be fairly easy... \[\frac{ 15 }{ 50 } = \frac{ x }{ 100 }\] multiply by 2!

OpenStudy (anonymous):

.03

OpenStudy (anonymous):

i mean .3

OpenStudy (anonymous):

15/50 equals .3 already

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=15%2F50

OpenStudy (anonymous):

50 is to 100 as 15 is to what?

OpenStudy (anonymous):

30??

OpenStudy (anonymous):

I'm so confused!!

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

Why are you confused?

OpenStudy (anonymous):

Idk I couldn't concentrate there for a second

OpenStudy (anonymous):

Look at it as a fraction! Here is an example \[\frac{ 1 }{ 2 } = \frac{ 2 }{ 4 } = \frac{ 20 }{ 40 }\]